# Unveiling Hallucination in Text, Image, Video, and Audio Foundation Models: A Comprehensive Review

### v1

@@ 2.3 Benchmark Evaluation
In certain instances, LLMs engage in a phenomenon termed "hallucination snowballing," where they fabricate false claims to rationalize prior hallucinations despite acknowledging their inaccuracy. To empirically explore this phenomenon, Zhang et al. (2023a) devise three question-answering datasets spanning diverse domains, wherein ChatGPT and GPT-4 often furnish inaccurate answers alongside explanations featuring at least one false claim. Significantly, the study suggested that the language model can discern these false claims as incorrect. Another benchmark dataset, FactCHD Chen et al. (2023b), was introduced to detect fact-conflicting hallucinations within intricate inferential contexts. It encompasses a range of datasets capturing different factual patterns and integrates fact-based evidence chains to improve assessment accuracy. Li et al. (2023b) introduced a dataset to assess the capability of LLMs to identify and recognize hallucinated or incorrect information. The outcomes highlighted ChatGPT’s inclination to produce hallucinated content, particularly on certain topics, introducing unverifiable information.
Figure 4: Example of types of Vision Hallucination.  Liu et al. (2024b) including visuals and the matching question-answer pairs and hallucination elements. While red annotated words do not exist or do not correspond to elements within the image, whereas words highlighted in green have correspondences within the image. Question and Answer are denoted by the letters Q and A, respectively.
3 Hallucination in Large Vision-Language Models
Large Vision-Language Models (LVLMs) have garnered significant attention in the AI community for their capacity to handle visual and textual data simultaneously. Nonetheless, similar to LLMs, LVLMs also confront the issue of hallucination. Figure 4 illustrates an example of visual hallucination.
@@ 3.1 Hallucination Detection and Mitigation
Dai et al. (2022) investigate the issue of object hallucinations in Vision-Language Pre-training (VLP) models, where textual descriptions generated by these models contain non-existent or inaccurate objects based on input images.  Li et al. (2023f) reveal widespread and severe object hallucination issues and suggest that visual instructions may influence hallucination. They observe that objects frequently depicted in visual instructions or co-occurring with image objects are more susceptible to hallucination. To enhance the evaluation process of object hallucination, the authors introduced a polling-based query method called POPE, which demonstrates improved stability and flexibility in assessing object hallucination. The absence of a standardized metric for assessing object hallucination has hindered progress in understanding and addressing this issue. To address this gap, Lovenia et al. (2023) introduced NOPE (Negative Object Presence Evaluation), a benchmark for evaluating object hallucination in vision-language models (VLMs) through visual question answering (VQA). Utilizing LLMs, the study generated a dataset of 29.5k synthetic negative pronoun (NegP) instances for NOPE. It thoroughly assessed the capability of 10 VLMs in detecting the absence of objects in visual questions, in addition to their typical performance on visual questions across nine other VQA datasets. Existing research focuses primarily on object hallucination, overlooking the LVLM hallucinations. Liu et al. (2024b) delve into Intrinsic Vision-Language Hallucination (IVL-Hallu) and proposes several novel IVL-Hallu tasks, including attribute, object, multi-modal conflicting, and counter-common-sense hallucination. They introduced a challenging benchmark dataset to assess and explore IVL-Hallu, conducting experiments on five LVLMs that revealed their limited effectiveness in addressing the proposed tasks. To mitigate object hallucination in LVLMs without resorting to costly training or API reliance, Zhao et al. (2024) introduced MARINE, a training-free and API-free solution. MARINE enhances LVLMs’ visual comprehension by combining existing open-source vision models and leveraging guidance without classifiers to incorporate object grounding features, thereby enhancing the precision of generated outputs. Evaluations across six LVLMs revealed MARINE’s effectiveness in reducing hallucinations and enhancing output detail, validated through assessments using GPT-4V.
HalluciDoctor Yu et al. (2023a) tackled hallucinations in Multi-modal Large Language Models (MLLMs) by using human error detection to identify and eliminate various types of hallucinations. Through rebalancing data distribution via counterfactual visual instruction expansion, they successfully mitigate 44.6% of hallucinations while maintaining competitive performance. Despite proficiency in visual semantic comprehension and meme humor, MLLMs struggle with chart analysis and understanding. Addressing this, Xu et al. (2023b) propose ChartBench, a benchmark assessing chart comprehension. ChartBench exposes MLLMs’ limited reasoning with complex charts, prompting the need for novel evaluation metrics like Acc+ and a handcrafted prompt, ChartCoT. Zhang et al. (2023b) introduced InternLM-XComposer, an LVLM aimed at designed to address the challenge of hallucination in image-text comprehension and composition. The performance of InternLM-XComposer’s text-image composition is evaluated through a robust procedure involving both human assessment and comparison to GPT4-Vision, with the model demonstrating competitive performance against solutions like GPT4-V and GPT3.5.
Cutting-edge LVLMs like InstructBLIP Dai et al. (2024), while generating visually grounded responses, often include inaccuracies like fictional items and flawed relationships. To enhance accuracy, Gunjal et al. (2024) introduced M-HalDetect, a pioneering multi-modal fine-grained hallucination detection dataset. This dataset serves as a benchmark for training LVLMs, leading to more precise outputs. Using fine-grained multi-modal reward models and enhancing FDPO significantly reduced InstructBLIP’s hallucination rate. These methods not only improve accuracy across LVLMs like LLaVA and mPLUG-OWL but also highlight the versatility and effectiveness of M-HalDetect in identifying and minimizing hallucinations.
Despite advancements in multi-modal tasks, LMMs often generate descriptions that are inconsistent with the accompanying image or human instructions. To address this, Liu et al. (2023) develop LRV-Instruction, a comprehensive dataset comprising 400k visual instructions across 16 tasks. This dataset includes both positive and negative instructions in various styles and semantic levels. Through LRV-Instruction, the issue of hallucination in existing LMMs was extensively examined, confirming its effectiveness in enhancing visual instruction tuning. Additionally, they introduced GAVIE, a novel method for evaluating visual instruction tuning without the need for human-labeled answers, which can be adapted to different types of instructions. The LVLM Hallucination Revisor (LURE) algorithm Zhou et al. (2023) was developed to correct object hallucination in LVLMs by refining the descriptions to produce more accurate and less hallucinatory outputs. Its methodology is based on an in-depth statistical analysis that identifies key factors contributing to object hallucination, such as the co-occurrence of certain objects in images, the uncertainty associated with objects during LVLM decoding, and the tendency for hallucinations to occur towards the end of the generated text. LURE is designed for seamless integration with a variety of LVLMs. When tested across multiple LVLMs, the integration of LURE resulted in significant enhancements in object hallucination correction, consistently outperforming other methods in both GPT and human evaluations based on various metrics.
@@ 3.2 Benchmark Evaluation
The current methods of developing LVLMs rely heavily on annotated benchmark datasets, which can exhibit domain bias and limit model generative capabilities. To address this, Li et al. (2023e) propose a novel data collection approach that synthesizes images and dialogues synchronously for visual instruction tuning, yielding a large dataset of image-dialogue pairs and multi-image instances.  Huang et al. (2024) introduced VHTest, a benchmark dataset with 1,200 diverse visual hallucinations (VH) instances across 8 VH modes. Evaluation of three SOTA MLLMs showed varying performance, with GPT-4V exhibiting lower hallucination than MiniGPT-v2. Rawte et al. (2024a) categorized visual hallucination in VLMs into eight orientations and introduced a dataset of 2,000 samples covering these types. They proposed three main categories of methods to mitigate hallucination: data-driven approaches, training adjustments, and post-processing techniques. Additionally, Wang et al. (2024) proposed the Visual Instruction Generation and Correction (VIGC) framework to address the shortage of high-quality instruction-tuning data for MLLMs. VIGC enables MLLMs to generate diverse instruction-tuning data while iteratively refining its quality through Visual Instruction Correction (VIC), mitigating hallucination risks. The framework produces diverse, high-quality data for fine-tuning models, validated through evaluations, improving benchmark performance, and overcoming language-only data limitations.
Figure 5: A video featuring descriptions generated by VLTinT model and ground truth (GT) with description errors highlighted in red italics Chuang and Fazli (2023).
4 Hallucinations in Large Video Models
Large Video Models (LVMs) represent a significant advancement, allowing for the processing of video data at scale. Despite their potential for various applications like video understanding and generation, LVMs face challenges with hallucinations, where misinterpretations of video frames can result in artificial or inaccurate visual data. This issue arises due to the video data complexity, which requires the model to process and comprehend it thoroughly. Figure 5 demonstrates the instances of hallucination observed in LVMs.
@@ 4.1 Hallucination Detection and Mitigation
The intricate task of dense video captioning, involving the creation of descriptions for multiple events within a continuous video, necessitates a thorough understanding of video content and contextual reasoning to ensure accurate description generation. However, this endeavor faces numerous challenges, potentially resulting in instances of inaccuracies and hallucinations Iashin and Rahtu (2020)Suin and Rajagopalan (2020). Traditional methods detect event proposals first, then caption subsets, risking hallucinations due to overlooking temporal dependencies. To address this, Mun et al. (2019) introduces a novel approach to modeling temporal dependencies and leveraging context for coherent storytelling. By integrating an event sequence generation network and a sequential video captioning network trained with reinforcement learning and two-level rewards, the model captures contextual information more effectively, yielding coherent and accurate captions while minimizing the risk of hallucinations.  Liu and Wan (2023) introduces a novel weakly-supervised, model-based factuality metric called FactVC, which outperforms previous metrics. Furthermore, they provide two annotated datasets to promote further research in assessing the factuality of video captions. Wu and Gao (2023) proposed a context-aware model that incorporates information from past and future events to influence the description of the current event conditionally. Their approach utilizes a robust pre-trained context encoder to encode information about the surrounding context events, which is then integrated into the captioning module using a gate-attention mechanism. Experimental findings on the YouCookII and ActivityNet datasets demonstrate that the proposed context-aware model outperforms existing context-aware and pre-trained models significantly. To enhance dense video captioning,  Zhou et al. (2024) introduces a streaming model comprising a memory module for long video handling and a streaming decoding algorithm enabling predictions before video completion. This approach notably boosts performance on prominent dense video captioning benchmarks, such as YouCook2, ActivityNet, and ViTT.
Video infilling and prediction tasks are crucial for assessing a model’s ability to comprehend and anticipate the temporal dynamics within video sequences Höppe et al. (2022). To address this, Himakunthala et al. (2023) introduced an inference-time challenge dataset containing keyframes with dense captions and structured scene descriptions. This dataset contains keyframes supplemented with unstructured dense captions and structured FAMOUS: (Focus, Action, Mood, Objects, and Setting) scene descriptions, providing valuable contextual information to support the models’ understanding of the video content. They employed various language models like GPT-3, GPT-4, and Vicuna with greedy decoding to mitigate hallucination risks.
Prominent developments in video inpainting have been observed recently, especially in situations where explicit guidance like optical flow helps to propagate missing pixels across frames Ouyang et al. (2021). However, difficulties and constraints occur from a lack of cross-frame information. Yu et al. (2023b) aims to tackle the opposite issue rather than depending on using pixels from other frames. The suggested method presents a Deficiency-aware Masked Transformer (DMT), a dual-modality-compatible inpainting framework. This approach improves handling scenarios with incomplete information by pre-training an image inpainting model to serve as a prior for training the video model.
Understanding scene affordances, which involve potential actions and interactions within a scene, is crucial for comprehending images and videos. Kulal et al. (2023) introduces a method for realistically inserting people into scenes. The model seamlessly integrates individuals into scenes by deducing realistic poses based on the context and ensuring visually pleasing compositions. Chuang and Fazli (2023) introduces CLearViD, a transformer-based model that utilizes curriculum learning techniques to enhance performance. By adopting this approach, the model acquires more robust and generalizable features. Furthermore, CLearViD incorporates the Mish activation function to address issues like vanishing gradients, thereby reducing the risk of hallucinations by introducing nonlinearity and non-monotonicity. Extensive experiments and ablation studies validate the effectiveness of CLearViD, with evaluations on ActivityNet Captions and YouCook2 datasets showcasing significant improvements over existing SOTA models in terms of diversity metrics.
@@ 4.2 Benchmark Evaluation
Zhang et al. (2006) created an innovative two-level hierarchical fusion method to hallucinate facial expression sequences from training video samples using only one frontal face image with a neutral expression. To effectively train the system, they introduced a dataset specifically designed for facial expression hallucination, which included 112 video sequences covering four types of facial expressions (happy, angry, surprise, and fear) from 28 individuals, resulting in the generation of reasonable facial expression sequences in both the temporal and spatial domains with less artifact. In the realm of video understanding, the development of end-to-end chat-centric systems has become a growing area of interest. Zhou et al. (2018) assembled the YouCook2 dataset, an extensive set of cooking videos with temporally localized and described procedural segments, to facilitate procedure learning tasks. Li et al. (2023c) introduced "VideoChat", a novel approach integrating video foundation models and LLMs through a learnable neural interface to enhance spatiotemporal reasoning, event localization, and causal relationship inference in video understanding. The researchers constructed a video-centric instruction dataset with detailed descriptions and conversations, emphasizing spatiotemporal reasoning and causal relationships. To counteract model hallucination, they employed a multi-step process to condense video descriptions into coherent narratives using GPT-4 and refined them to improve clarity and coherence. To explore the challenge of deducing scene affordances, Kulal et al. (2023) curated a dataset of 2.4M video clips, showcasing a variety of plausible poses that align with the scene context.
Figure 6: Audio hallucination examples for each classes - Type A: Involving hallucinations of both objects and actions $|$ Type B: Featuring accurate objects but hallucinated actions $|$ Type C: Displaying correct actions but hallucinated objects Nishimura et al. (2024).
5 Hallucinations in Large Audio Models
Large audio models (LAMs) have emerged as a powerful tool in the realm of audio processing and generation, with a wide range of applications like speech recognition, music analysis, audio synthesis, and captioning Latif et al. (2023)Ghosal et al. (2023). Although these models have demonstrated remarkable capabilities across various domains, they are susceptible to hallucinations. These anomalies can take several forms, from creating unrealistic audio by piecing together fabricated snippets to injecting false information, such as quotes or facts, into summaries. Additionally, they may fail to accurately capture the inherent features of audio signals, such as timbre, pitch, or background noise Shen et al. (2023).
5.1 Hallucination Detection and Mitigation
In the realm of audio captioning, where natural language descriptions for audio clips are automatically generated, a significant challenge arises from the over-reliance on the visual modality during the pre-training of audio-text models. This reliance introduced data noise and hallucinations, ultimately undermining the accuracy of the resulting captions. To address this issue, Xu et al. (2023a) introduces an AudioSet tag-guided model designed to bootstrap large-scale audio-text data (BLAT). Notably, this model sidesteps the incorporation of video, thus minimizing noise associated with the visual modality. The experimental findings across a range of tasks, including retrieval, generation, and classification, validate the effectiveness of BLAT in mitigating hallucination issues.
Speech emotions play a crucial role in human communication and find extensive applications in areas such as speech synthesis and natural language understanding. However, traditional categorization approaches may fall short of capturing the nuanced and intricate nature of emotions conveyed in human speech Jiang et al. (2019)Han et al. (2021)Ye et al. (2021). SECap Xu et al. (2024a), a framework designed for speech emotion captioning, aims to capture the intricate emotional nuances of speech using natural language. SECap utilizes various components, including LLaMA as the text decoder, HuBERT as the audio encoder, and Q-Former as the Bridge-Net, to generate coherent emotion captions based on speech features. Audio-language models, despite their capability for zero-shot inference, confront challenges like hallucinating task-specific details despite strong performance. To address this, Elizalde et al. (2024) introduces the Contrastive Language-Audio Pretraining (CLAP) model. Pre-trained with 4.6 million diverse audio-text pairs, CLAP features a dual-encoder architecture, enhancing representation learning for improved task generalization across sound, music, and speech domains.
5.2 Benchmark Evaluation
@@
To address the scarcity of data in the specific domain of music captioning, Doh et al. (2023) introduced LP-MusicCaps, a comprehensive dataset comprising 0.5 million audio clips accompanied by approximately 2.2 million captions. Leveraging LLMs, they train a transformer-based music captioning model with the dataset and assess its performance under zero-shot and transfer-learning scenarios, demonstrating its superiority over supervised baseline models.  Nishimura et al. (2024) investigated audio hallucinations in large audio-video language models, where audio descriptions are generated primarily based on visual information, neglecting audio content. They have classified these hallucinations into three distinct types such as Involving hallucinations of both objects and actions, Featuring accurate objects but hallucinated actions, and Displaying correct actions but hallucinated objects as represented in Fig. 6. In their investigation, they gathered 1000 sentences by soliciting audio information and then annotated them to determine whether they contained auditory hallucinations, further categorizing the type of hallucination if detected. To assess compositional reasoning in LAMs, Ghosh et al. (2023) introduced CompA, consisting of two expert-annotated benchmarks primarily focused on real-world audio samples. This benchmark is employed to fine-tune CompA-CLAP with a novel learning approach, enhancing its compositional reasoning skills and demonstrating substantial improvement over all the baseline models in tasks requiring compositional reasoning.
TEXT Paper Detection Mitigation Task Dataset(s) Evaluation Metric(s)
Manakul et al. (2023) Yes No QA Wikibio Entropy
Li et al. (2022) Yes Yes QA, Dialog summarization Halueval Automatic
Mündler et al. Yes Yes Text generation Manual F1 Score
Chen et al. (2023a) No Yes Editing for attribution MCQ, Dialog Attribution, Preservation
Zhang et al. (2023c) No Yes Question knowledge alignment Fuzzy QA Attributable to Identified Sources
Zhang et al. (2023a) Yes No QA Manual Accuracy
Peng et al. (2023) No Yes Task-oriented dialog News, Customer service F1 Score, Bleu-4
Cui et al. (2023) No Yes QA Manual Ranking
Azaria and Mitchell (2023) Yes No Classification Manual Accuracy
Li et al. (2023d) Yes Yes Knowledge-intensive tasks Fever, QA Accuracy
Elaraby et al. (2023) Yes Yes Consistency, Actuality, QA Manual NBA domain Pearson Correlation Coefficient
Varshney et al. Yes Yes Text generation Wikibio Percentage of mitigated hallucination
Jha et al. (2023) Yes No Dialog N/A N/A
Pal et al. (2023) No No Reasoning hallucination Med-Halt Accuracy, Pointwise Score
McKenna et al. (2023) Yes No Textual entailment Altered Directional Inference Entailment Probability
Guerreiro et al. (2023) Yes Yes MT FLores 101, WMT ,TICO BLEU
Huang and Chang (2023) Yes Yes N/A N/A N/A
Luo et al. (2023) Yes Yes Concept extraction Concept-7 AUC, Accuracy, F1 Score
Gao et al. (2022) Yes Yes Editing attribution NQ, SQA Auto-AIS (Attr_auto)
Yang et al. (2023) Yes No
Detect factual
errors automatically
PHD, WikiBio-GPT3
Precision, Recall,
F1 Score, Accuracy
Min et al. (2023) Yes Yes Fact verification Manual(Wikipedia) FActScore
Rawte et al. (2024b) Yes Yes Factual inaccuracies detection FACTOID HV I_auto
Ahmad et al. (2023) Yes Yes Hallucination in healthcare N/A FActScores
Ji et al. (2023) Yes Yes Generative and knowledge-intensive
PubMedQA, MEDIQA2019,
MedQuAD, and MASH-QA
Unigram F1, ROUGE-L,
Med-NLI, and CTRLEval
Kang and Liu (2023) Yes Yes Hallucination in finance N/A FActScores
Roychowdhury (2024) No Yes QA N/A N/A
Savelka et al. (2023) No Yes Factual evaluation in legislation N/A N/A
Dahl et al. (2024) Yes No Legal hallucination Manual N/A
Rawte et al. (2024c) No Yes Comprehension enhancement SCA-90K Cosine similarity
IMAGE Li et al. (2023f) Yes No Evaluation of object hallucination MSCOCO CHAIR, POPE
Gunjal et al. (2024) Yes Yes VQA M-Hall Detect Accuracy
Dai et al. (2022) No Yes Image captioning CHAIR CIDEr
Lovenia et al. (2023) Yes No Object hallucination NOPE
METEOR, Exact match accuracy,
NegP Accuracy
Liu et al. (2024b) Yes No Intrinsic vision-language hallucination PhD Accuracy
Zhao et al. (2024) Yes Yes Non-existing object hallucination MSCOCO CHAIR, POPE, GPT-4V, recall
Huang et al. (2024) Yes No Visual hallucination YNQ, OEQ Accuracy
Rawte et al. (2024a) Yes No Video captioning ActivityNet-Fact, YouCook2-Fact FactVC
Wang et al. (2024) No Yes
Generate instruction data
for vision-language
VIGC-LLaVA-COCO,
VIGC-LLaVA-Objects365
Conv, Detail, Complex
Yu et al. (2023a) Yes Yes
Machine-generated
visual instruction
LLaVA-Instruction-158K CHAIR
Guan et al. (2023) No Yes Visual questions HallusionBench Accuracy
Liu et al. (2023) Yes Yes Vision language LRV-Instruction GAVIE
Xu et al. (2023b) Yes No
Evaluation of MLLMs on
chart comprehension
ChartBench Acc+
Lu et al. (2024) Yes Yes Vision language MSG-MCQ Accuracy
Tong et al. (2024) Yes No Visual question answering MMVP, VQA Accuracy
Liao et al. (2023) Yes No Vision language REVO-LION Meta Quality (MQ)
Hu et al. (2023) Yes Yes
Visual captioning,
Visual question answering
CIEM
Accuracy, Precision,
Recall, F1 Score
Jing et al. (2023) Yes No Meta-evaluation LLaVA-1k, MSCOCO-Cap FAITHSCORE
Changpinyo et al. (2022) No Yes Multilingual visual question answering MaXM Accuracy
Wang et al. (2023) Yes No Content generation N/A Precision, Recall, F1 Score
Sun et al. (2023) No Yes Visual-language alignment MMHAL-BENCH N/A
Bai et al. (2023) Yes No
Evaluate hallucination of
vision language model
TouchStone Hallucination Score
Zhou et al. (2023) No Yes Hallucination mitigation in LVMs MSCOCO CHAIR, BLEU, CLIP
Yan et al. (2024) No Yes Visual grounding MMViG
HL, CA, AA,
RA, RL, RS, DL
Zhao et al. (2023) Yes Yes Overcome hallucination in LVMs POPE, SHR Accuracy, Precision, F1 Score
Zhang et al. (2023b) No Yes
Image text comprehension
and composition
MMBench, SeedBench, QBench,
MMBench-CN, Chinese Bench
LR, AR, $R$,
FP-C, FP-S, CP
VIDEO Kulal et al. (2023) No Yes Affordance prediction Manual FID, FCKh
Himakunthala et al. (2023) No Yes Video infilling, Scene prediction Manual N/A
Li et al. (2023c) No Yes Visual dialogue Manual N/A
Zhou et al. (2024) No Yes Video captioning
ActivityNet Captions,
YouCook2, ViTT
CIDER, METEOR, SODAc
Höppe et al. (2022) Yes No Video prediction BAIR, Kinetics 600, UCF-101 Frechet Video Distance
Chuang and Fazli (2023) No Yes Video description Activity Net Captions, YouCook2
METEOR, ROUGE_L, CIDER,
BLEU_4, DIV-2, RE_4
AUDIO Li et al. (2023a) No Yes Classification Manual Mean avg precision
Doh et al. (2023) No Yes Audio captioning LP MusicCaps BLEU
Xu et al. (2023a) No Yes Caption generation AudioCaps R@K, COCO & FENCE
Liu and Wan (2023) No Yes Audio captioning MusciCaps BLEU
Nishimura et al. (2024) Yes No Evaluation of LAMs LAION_CLAP,MS_CLAP Recall, Precision, F1 Score
Table 1: Overview of the hallucination detection and mitigation landscape in FMs across modalities (Text, Image, Video, and Audio). Each work is categorized based on factors such as detection, mitigation, tasks, datasets, and evaluation metrics.
6 Hallucination: Good or Bad?
Hallucinations in large-scale models present a complex interplay between creativity and uncertainty. On one hand, the ability to traverse beyond conventional data boundaries can lead to the generation of novel and innovative outputs. Hallucinations can spark exploratory learning, revealing unexpected patterns and features within the data. They can also serve as a form of stress testing, improving the model’s robustness and adaptability. Furthermore, these unexpected outputs can even inspire human creativity, serving as a springboard for new ideas and perspectives Rawte et al. (2023b). However, this dual nature of hallucinations also introduced significant drawbacks. The quality and coherence of hallucinatory outputs can be questionable, posing challenges in applications where accuracy and reliability are paramount. Hallucinations can also propagate misinformation and biases present in the model’s training data, potentially reinforcing existing prejudices and eroding user trust. The reduced interpretability of these outputs can further undermine the model’s credibility and adoption. Ethical concerns arise when hallucinations produce inappropriate, offensive, or harmful content. Careful monitoring and control mechanisms are essential to prevent the generation of outputs that could cause harm or distress to users. Navigating this intricate balance between exploration and fidelity is crucial for maximizing the utility of large models while mitigating the risks associated with unexpected outputs. Overall, the phenomenon of hallucinations in large-scale models highlights the need for a nuanced understanding and strategic management of these capabilities.
7 Limitations
Previous survey papers primarily focused on hallucination in Large Language Models and did not extensively cover hallucinations in vision, audio, and video modalities. In this survey paper, our aim is to provide a comprehensive overview of hallucinations across all modalities, considering that hallucinations can occur in any large foundation model. Despite our efforts to provide a comprehensive summary of recent advancements related to hallucination techniques in all foundational models, we acknowledge that we may miss some relevant work in the field.

### v4

@@ 3.2 Benchmark Evaluation
The current methods of developing LVLMs rely heavily on annotated benchmark datasets, which can exhibit domain bias and limit model generative capabilities. To address this, Li et al. (2023f) proposed a novel data collection approach that synthesizes images and dialogues synchronously for visual instruction tuning, yielding a large dataset of image-dialogue pairs and multi-image instances.  Huang et al. (2024b) introduced VHTest, a benchmark dataset with 1,200 diverse visual hallucinations (VH) instances across 8 VH modes. Evaluation of three SOTA MLLMs showed varying performance, with GPT-4V exhibiting lower hallucination than MiniGPT-v2.  Rawte et al. (2024a) categorized visual hallucination in VLMs into eight orientations and introduced a dataset of 2,000 samples covering these types. They proposed three main categories of methods to mitigate hallucination: data-driven approaches, training adjustments, and post-processing techniques. Wang et al. (2024a) proposed the Visual Instruction Generation and Correction (VIGC) framework to address the scarcity of high-quality instruction-tuning data for MLLMs. VIGC enables MLLMs to generate diverse instruction-tuning data while iteratively refining its quality through Visual Instruction Correction (VIC), mitigating hallucination risks. The framework produces diverse, high-quality data for fine-tuning models, validated through evaluations, improving benchmark performance, and overcoming language-only data limitations.
Figure 5: A video featuring descriptions generated by VLTinT model and ground truth (GT) with description errors highlighted in red italics. Chuang and Fazli (2023).
4 Hallucinations in Large Video Models
Large video models (LVMs) represent a significant advancement, allowing for processing video data at scale. Despite their potential for various applications like video understanding and generation, LVMs face challenges with hallucinations, where misinterpretations of video frames can result in artificial or inaccurate visual data. Figure 5 demonstrates the instances of hallucination observed in LVMs.
@@ 4.1 Hallucination Detection and Mitigation
The intricate task of dense video captioning, involving the creation of descriptions for multiple events within a continuous video, necessitates a thorough understanding of video content and contextual reasoning to ensure accurate description generation. However, this endeavor faces numerous challenges, potentially resulting in instances of inaccuracies and hallucinations Iashin and Rahtu (2020)Suin and Rajagopalan (2020). Traditional methods detect event proposals first, then caption subsets, risking hallucinations due to overlooking temporal dependencies. To address this, Mun et al. (2019) introduced a novel approach to modeling temporal dependencies and leveraging context for coherent storytelling. By integrating an event sequence generation network and a sequential video captioning network trained with reinforcement learning and two-level rewards, the model captures contextual information more effectively, yielding coherent and accurate captions while minimizing the risk of hallucinations. Liu and Wan (2023) introduced a novel weakly-supervised, model-based factuality metric called FactVC, which outperforms previous metrics. Furthermore, they provided two annotated datasets to promote further research in assessing the factuality of video captions. Wu and Gao (2023) proposed a context-aware model that incorporates information from past and future events to influence the description of the current event conditionally. Their approach utilizes a robust pre-trained context encoder to encode information about the surrounding context events, which is then integrated into the captioning module using a gate-attention mechanism. Experimental findings on the YouCookII and ActivityNet datasets demonstrate that the proposed context-aware model outperforms existing context-aware and pre-trained models by a significant margin. To enhance dense video captioning,  Zhou et al. (2024) introduced a streaming model comprising a memory module for long video handling and a streaming decoding algorithm enabling predictions before video completion. This approach notably boosts performance on dense video captioning benchmarks such as ActivityNet, YouCook2, and ViTT.
Video infilling and prediction tasks are crucial for assessing a model’s ability to comprehend and anticipate the temporal dynamics within video sequences Höppe et al. (2022). To address this, Himakunthala et al. (2023) introduced an inference-time challenge dataset containing keyframes with dense captions and structured scene descriptions. This dataset contains keyframes supplemented with unstructured dense captions and structured FAMOUS: (Focus, Action, Mood, Objects, and Setting) scene descriptions, providing valuable contextual information to support the models’ understanding of the video content. They employed various language models like GPT-3, GPT-4, and Vicuna with greedy decoding to mitigate hallucination risks. Prominent developments in video inpainting have been observed recently, especially in situations where explicit guidance like optical flow helps to propagate missing pixels across frames Ouyang et al. (2021). However, difficulties and constraints occur from a lack of cross-frame information. Yu et al. (2023b) aimed to tackle the opposite issue rather than depending on using pixels from other frames. The suggested method presents a Deficiency-aware Masked Transformer (DMT), a dual-modality-compatible inpainting framework. This approach improves handling scenarios with incomplete information by pre-training an image inpainting model to serve as a prior for training the video model.
Understanding scene affordances, which involve potential actions and interactions within a scene, is crucial for comprehending images and videos. Kulal et al. (2023) introduced a method for realistically inserting people into scenes. The model seamlessly integrates individuals into scenes by deducing realistic poses based on the context and ensuring visually pleasing compositions.  Chuang and Fazli (2023) introduced CLearViD, a transformer-based model that utilizes curriculum learning techniques to enhance performance. By adopting this approach, the model acquires more robust and generalizable features. Furthermore, CLearViD incorporates the Mish activation function to address issues like vanishing gradients, thereby reducing the risk of hallucinations by introducing nonlinearity and non-monotonicity. Extensive experiments and ablation studies validate the effectiveness of CLearViD, with evaluations on ActivityNet captions and YouCook2 datasets showcasing significant improvements over existing SOTA models in terms of diversity metrics.
@@ 4.2 Benchmark Evaluation
Zhang et al. (2006) created a novel two-level hierarchical fusion method to hallucinate facial expression sequences from training video samples using only one frontal face image with a neutral expression. To effectively train the system, they introduced a dataset specifically designed for facial expression hallucination, which included 112 video sequences covering four types of facial expressions (happy, angry, surprise, and fear) from 28 individuals, resulting in the generation of reasonable facial expression sequences in both the temporal and spatial domains with less artifact. In the realm of video understanding, the development of end-to-end chat-centric systems has become a growing area of interest. Zhou et al. (2018) assembled the YouCook2 dataset, an extensive set of cooking videos with temporally localized and described procedural segments, to facilitate procedure learning tasks. Li et al. (2023d) introduced "VideoChat," a novel approach integrating video foundation models and LLMs through a learnable neural interface to enhance spatiotemporal reasoning, event localization, and causal relationship inference in video understanding. The researchers constructed a video-centric instruction dataset with detailed descriptions and conversations, emphasizing spatiotemporal reasoning and causal relationships. To counteract model hallucination, they employed a multi-step process to condense video descriptions into coherent narratives using GPT-4 and refined them to improve clarity and coherence. To explore the challenge of deducing scene affordances, Kulal et al. (2023) curated a dataset of 2.4M video clips, showcasing a variety of plausible poses that align with the scene context.
Figure 6: Audio hallucination examples for each classes - Type A: Involving hallucinations of both objects and actions: Type B: Featuring accurate objects but hallucinated actions; Type C: Displaying correct actions but hallucinated objects Nishimura et al. (2024).
5 Hallucinations in Large Audio Models
Large audio models (LAMs) have emerged as a powerful tool in the realm of audio processing and generation, with a wide range of applications like speech recognition, music analysis, audio synthesis, and captioning Latif et al. (2023)Ghosal et al. (2023). While demonstrating remarkable capabilities, LAMs are susceptible to hallucinations – anomalies ranging from generating unrealistic audio by combining fabricated snippets to injecting false information like quotes or facts into summaries. Additionally, they may fail to accurately capture the inherent features of audio signals, such as timbre, pitch, or background noise Shen et al. (2023). Figure 6 presents one example of audio hallucinations.
@@ 5.1 Hallucination Detection and Mitigation
In the realm of audio captioning, where natural language descriptions for audio clips are automatically generated, a significant challenge arises from the over-reliance on the visual modality during the pre-training of audio-text models. This reliance introduces data noise and hallucinations, ultimately undermining the accuracy of the resulting captions. To address this issue, Xu et al. (2023a) introduced an AudioSet tag-guided model designed to bootstrap large-scale audio-text data (BLAT). Notably, this model sidesteps the incorporation of video, thus minimizing noise associated with the visual modality. The experimental findings across a range of tasks, including retrieval, generation, and classification, validate the effectiveness of BLAT in mitigating hallucination issues.
Speech emotions play a crucial role in human communication and find extensive applications in areas such as speech synthesis and natural language understanding. However, traditional categorization approaches may fall short of capturing the nuanced and intricate nature of emotions conveyed in human speech Jiang et al. (2019)Han et al. (2021)Ye et al. (2021). SECap Xu et al. (2024a), a framework designed for speech emotion captioning. It aims to capture the intricate emotional nuances of speech using natural language. SECap utilizes various components, including LLaMA as the text decoder, HuBERT as the audio encoder, and Q-Former as the Bridge-Net, to generate coherent emotion captions based on speech features. Audio-language models, despite their capability for zero-shot inference, confront challenges like hallucinating task-specific details despite strong performance. To address this, Elizalde et al. (2024) introduced the Contrastive Language-Audio Pretraining (CLAP) model. Pre-trained with 4.6 million diverse audio-text pairs, CLAP features a dual-encoder architecture, enhancing representation learning for improved task generalization across sound, music, and speech domains.
@@ 5.2 Benchmark Evaluation
To address the scarcity of data in the specific domain of music captioning, Doh et al. (2023) introduced LP-MusicCaps, a comprehensive dataset comprising 0.5 million audio clips accompanied by approximately 2.2 million captions. Leveraging LLMs, they trained a transformer-based music captioning model with the dataset and assessed its performance under zero-shot and transfer-learning scenarios, demonstrating its superiority over supervised baseline models. Nishimura et al. (2024) investigated audio hallucinations in large audio-video language models, where audio descriptions are generated primarily based on visual information, neglecting audio content. They have classified these hallucinations into three distinct types such as Involving hallucinations of both objects and actions, Featuring accurate objects but hallucinated actions, and Displaying correct actions but hallucinated objects as represented in Fig. 6. They gathered 1000 sentences by soliciting audio information and then annotated them to determine whether they contained auditory hallucinations. To assess compositional reasoning in LAMs, Ghosh et al. (2023) introduced CompA, consisting of two expert-annotated benchmarks and employed to fine-tune CompA-CLAP with a novel learning approach, leading to improved compositional reasoning abilities compared to baseline models on related tasks.